In 2018, Simon, age 12, has interest income of $8,960 and no earned income. He has no investment expenses.

Simon's total tax is:
Answer #1

Taxable income

Interest income

8960

Less: standard deduction

(1050)

Taxable income

7910

Net unearned income

Interest income

8960

Less: standard deduction

(1050)

Less: statutory deduction

(1050)

Net unearned income

6860

Earned taxable income

Taxable income

7910

Net unearned income

(6860)

Earned taxable income

1050

First, Earned income + 2550, 1050+2550 = 3600 will be taxed at 10% and tax on it will be $360 (3600*10%)

Next remaining 4310 (7910-3600) will be taxed at 24% and tax on it will be 1034.40 (4310*24%)

Therefore,

Simon's total tax = 360+1034.40 = $1394.40